OpenStudy (anonymous):

3xsqrt of(x-2) = 2x sqrt of x

OpenStudy (anonymous):

\[3x \sqrt{(x-2)} = 2x \sqrt{x}\] Square both sides to get 9x^2(x-2)=4x^3 Distribute to get 9x^3 - 18x^2 = 4x^3 Subtract 4x^3 to get 5x^3 -18x^2 = 0 Factor x^2 out of both terms to get x^2(5x-18) Set each of the terms (x,x, and 5x-18) equal to zero to get: x=0, or 18/5